OpenStudy (tiff2002):

Calculate the monthly finance charge for the following credit card transaction. Assume that it takes 10 days for a payment to be received and recorded and that the month is 30 days long. (Round your answer to the nearest cent.) $500 balance, 20% rate, $50 payment, previous balance method

OpenStudy (kropot72):

Monthlt charge is as follows: \[(500\times \frac{0.2}{36})+(450\times \frac{0.4}{36})\]